# TAPIP3D: Tracking Any Point in Persistent 3D Geometry

### v2

@@ 1 Introduction
Tracking points over time in video—especially through occlusions—has become a valuable tool in robotics and action recognition [18, 17, 12]. Fine-grained motion estimation at the particle level offers a unified framework for capturing temporal changes in object pose, part articulation, deformable structures, and even granular materials. However, most existing point trackers operate either directly in pixel space [18, 17, 12] or in pixel space augmented with depth information [44, 28]. Yet, the predominant source of apparent motion in video is often camera movement, not object motion. Since real-world dynamics unfold in 3D, tracking points directly in 3D space—rather than in the image plane—may be more natural and effective.
@@
We test TAPIP3D in established 3D point tracking benchmarks of TAPVid3D [19], LSFOdyssey [37], Dynamic Replica [16] and DexYCB [4] for tracking points in 2D and 3D, using depth from ground-truth, sensors, or estimators, in camera and world 3D coordinate frames. We show TAPIP3D outperforms all previous methods in 3D point tracking metrics, with an especially large margin when accurate depth is available. While existing methods pursue 3D tracking in camera coordinates [28, 44], our experiments show that it is better to track in a “world” coordinate frame, which can be estimated by recent camera and depth estimation methods such as MegaSAM [23]. In addition to showing the impact of coordinate system choices, we ablate the contributions to performance from 3D-centric featurization and quality of depth estimation.
In summary, we present TAPIP3D for 3D point tracking from RGB-D and RGB videos that uses 3D feature clouds for video to estimate 3D point tracks, which demonstrates state-of-the-art performance when ground-truth / sensor depth is available and competitive performance with estimated depth [23]. It can exploit recent advances in depth and camera pose estimation [23, 43] and deliver both camera and world-centric 3D tracks. To our knowledge, TAPIP3D is a pioneering 3D tracking method capable of tracking in a 3D world-centric space, with camera motion factored out. We will release our code upon acceptance.
@@ 2 Related Work
Recent advances in point tracking formulate the task as the estimation of multi-frame point trajectories, moving beyond traditional pairwise optical flow estimation. Inspired by earlier work on particle video [33], PIPs [12] introduces an iterative refinement approach with dense cost maps. This work along with the Tracking Any Point (TAP) benchmark [7] have catalyzed a series of advancements [21, 22] including the use of multi-point context [18, 17], enhanced initialization schemes [8], broader correlation contexts [2, 5], and post-hoc densification techniques [20].
Beyond 2D, recent works such as SpatialTracker [44] have extended point tracking to 3D by incorporating depth information. SpatialTracker employs a Triplane representation [3], where 3D featured point clouds per frame from pixel (UV) and depth (D) coordinates are projected onto onto three orthogonal planes. Triplane featurization is faster than our -NN attention, but this speedup comes at the cost of decreased performance. The works of DELTA [28] and SceneTracker [37] also operate on a UVD coordinate system to separately compute appearance and depth correlation features, directly extending the 2D method of CoTracker [18]. DELTA [28] proposes a coarse-to-fine trajectory estimation that manages to estimate dense point tracks across the whole image plane, instead of on a set of sparse locations. In contrast to previous 3D point trackers, TAPIP3D leverages an explicit spatio-temporal 3D feature cloud video representation, instead of a 2.5D one of the image-space and depth, for feature extraction and tracking. By fully utilizing the underlying geometric structure of the scene, our method achieves better point tracking accuracy and consistency. Moreover, our model can estimate trajectories in a 3D world space, by canceling inferred camera motion from the scene image pixels during lifting, which helps performance further.
@@
3D point tracks can also be extracted from monocular videos through test-time optimization by fitting dynamic radiance fields, such as dynamic NeRFs [31, 38] or Gaussians [27, 39, 34, 6] to the observed video frames supervised from reprojection RGB, depth and motion error. These approaches require per-video optimization and are thus computationally expensive. In contrast, TAPIP3D is a learning-based feed-forward approach for 3D point tracking.
@@
Scene flow estimation extends optical flow into three dimensions by considering pairs of point clouds and estimating the 3D motion between them [11, 13, 32, 36, 25, 42, 10, 29]. Recent works incorporate rigid motion priors, explicitly [36] or implicitly [45] or leveraging diffusion models [24]. Linking 3D scene flow estimates terminates at pixel occlusions by design, similarly to linking 2D flow estimates. Instead our work focuses on inferring multiframe point trajectories through occlusions.
@@
Recent methods in learning based camera motion and depth estimation, such as MoGe [40], DUSt3R [41], MonST3R [46], and MegaSaM [23], focus on predicting dense 3D reconstructions from single images, image pairs, or videos, bringing the vision of video to 4D translation closer. MegaSAM [23] delivers highly accurate camera motion and depth by combining learning based initialization and updates with second order optimization, extending earlier work of DROID-SLAM [35]. None of the methods above addresses the problem of estimating the 3D point motion, which is the focus of our work. TAPIP3D builds upon the progress of recent learning-based 3D camera motion and depth estimation methods to explore their use in 3D point tracking, by lifting videos to a 3D world space which makes the scene points easier to track than in the original 2D image plane.
@@ 3 Method
Concretely, we take as input an RGB video , a sequence of depth maps , camera intrinsics (constant over time), and optionally, a sequence of camera poses. The input depth maps can originate from an off-the-shelf monocular depth estimator [40], depth sensors, or GT depth provided by simulation environments. Camera intrinsics and extrinsics can either be estimated algorithmically by video 3D reconstruction models [23] or be provided by the dataset. We are given a set of different query points , , where represents the timestep at which a query is provided. Queries may also be specified in 2D, in which case we lift them to 3D using depth and camera intrinsics before sending them to the model. For each query, our model produces a 3D point trajectory , and a visibility trajectory . The architecture of TAPIP3D is illustrated in Figure 2.
@@
Previous 3D point trackers [28, 37, 44] represent 3D point trajectories by concatenating depth values to pixel coordinates, which we refer to as space. In contrast, we use coordinates, which allows us to better leverage 3D priors, and also compensate for camera motion by transforming all frames’ points into the first camera’s coordinate system (if poses are available). Disentangling camera motion from scene motion is helpful because it makes the 3D tracks more predictable.
@@ 3.1 Preliminaries: Point tracking as iterative trajectory refinement
TAPIP3D builds on the trend of iterative architectures for point tracking [12, 8, 18, 17]. We build from the most recent such tracker, CoTracker3 [17].
@@
CoTracker3 tracks 2D points across time by iteratively refining their coordinates and visibilities through a transformer. Given a sequence of frames and a set of query points, for each query it maintains and updates per-frame (2D) locations , and confidences , and visibilities . The trajectory is initialized by copying the query coordinate to all frames (zero motion assumption); and are initialized with zeros. The model then performs iterations of refinement, producing additive updates to the coordinates and visibilities and confidences.
@@ 3.2 Tracking in a Persistent 3D Geometric Space
Our key innovation is to represent the input video as a sequence of multi-scale 2D feature maps enriched with 3D coordinates: each 2D cell holds both a -dimensional feature vector and a 3D position, where the 3D positions come from unprojected depths. This representation allows us to treat the data as a sequence of featurized point clouds (or “feature clouds”), despite their 2D grid arrangement in memory. Compared to other 3D scene representations such as tri-plane encoders [44] and UVD coordinates [28], our representation encodes 3D geometry more explicitly and without distortions or approximations. We first encode each video frame with a 2D image encoder with downsampling rate , producing a sequence of feature maps at a reduced resolution , where and are the original width and height of the image. We convert these into levels of multi-scale features via 2D downsampling: for . Our downsampling uses nearest-neighbor interpolation for coordinates and average pooling for features.
@@
Our next step is to contextualize each timestep of each trajectory () with information local to it in the input data. Previous methods compute correlation features for each trajectory point using 2D neighborhoods, either on the image plane [18, 5] or on triplanes [44]. Our representation enables us to use 3D neighborhoods at this step, which lets the model take advantage of the fact that the targets are in a 3D world and are often disambiguated by 3D distance. Our overall strategy is illustrated in Figure 3.
@@
These trajectory tokens are processed by a transformer utilizing proxy tokens, following CoTracker3 [17]. That is, we use virtual trajectories (that do no correspond to any query points) to help approximate spatial and temporal attention across our main tokens. The transformer outputs additive updates for positions and visibility scores:
@@
Our model is trained on the Kubric MOVi-F dataset [9]. We initialize the image encoder with CoTracker3’s pre-trained weights [17]. We train on 8 L40S GPUs with a batch size of 1 for 200K iterations, which takes roughly 4.2 days and consumes only 20GB VRAM with BF16 mixed precision. We train and test with a window length of . We train with videos of length 24 (making two windows per batch), and trajectories per sample. We scale the coordinates in the point maps to unit variance before processing each window to simplify learning. We optimize using AdamW [26] with the learning rate and weight decay both set to 5e-4. We use a learning rate schedule with 10k warmup steps followed by cosine annealing. During inference, under BF16 mixed precision, our model achieves a speed of 10 FPS and consumes around 2.6GB of VRAM when tracking 1024 query points across 32 frames on an L40S GPU.
@@ 4 Experiments
We adopt the 3D point tracking metrics from TAPVid-3D [19]: APD3D () quantifies the average percentage of points whose errors fall within the threshold . Occlusion Accuracy (OA) assesses the precision of occlusion predictions, while Average Jaccard (AJ) jointly measures accuracy in both position and occlusion estimation.
@@
| CoTracker3 [17] + M-SaM | 20.4 | 30.1 | 89.8 | 14.1 | 20.3 | 88.5 | 17.4 | 27.2 | 85.0 | 17.3 | 25.9 | 87.8 |
| SpatialTracker [44] + M-SaM | 15.9 | 23.8 | 90.1 | 7.7 | 13.5 | 85.2 | 15.3 | 25.2 | 78.1 | 13.0 | 20.8 | 84.5 |
| DELTA [28] + M-SaM | 21.0 | 29.3 | 89.7 | 14.6 | 22.2 | 88.1 | 17.7 | 27.3 | 81.4 | 17.8 | 22.3 | 86.4 |
@@
We evaluate the performance of our model and baselines in the following synthetic and real world datasets: 1) TAPVid3D [19] integrates videos from three distinct real-world datasets covering diverse scenarios: DriveTrack [1], Panoptic Studio [15], and Aria Digital Twin [30]. Together, these datasets provide a total of 4,569 evaluation videos, with video lengths ranging from 25 to 300 frames. 2) DexYCB-Pt: a dataset we introduce by leveraging ground-truth object and hand poses from manipulation scenes in DexYCB [4] to generate accurate 3D point tracks on 8,000 real-world RGB-D videos. 3) LSFOdyssey [37]: a synthetic dataset that contains 90 videos with complex but realistic long-range motion using humanoids, robots, and animals. Each video is 40 frames long and has 3D point trajectories and occlusion labels, sourced from PointOdyssey [47]. 4) DynamicReplica [16]: a synthetic 3D dataset that contains 500 videos of articulated human and animal models. Each video is 300 frames long.
@@
We compare TAPIP3D against the following 2D and 3D point trackers: 1) DELTA [28]: a very recent state-of-the-art 3D point tracker that uses depth as an additional input for computing cross-correlations, essentially extending CoTracker [18] to a UVD space. 2) SpatialTracker [44]: a 3D point tracker that builds upon CoTracker’s architecture and uses a tri-plane representation to featurize coordinates. 3) CoTracker3 [17], which operates in 2D, but we lift its trajectories to 3D using the depth at the trajectory coordinates.
@@
In all the evaluations, our model is trained on the Kubric MOVi-F dataset [9]. TAPIP3D-world and TAPIP3D-camera adopt the same trained checkpoint but conduct inference in the two different 3D coordinate systems.
@@ 4.1 Point Tracking Evaluation
We first evaluate TAPIP3D against strong baseline methods [17, 44, 28] on the challenging real-world TAPVid-3D [19] benchmark. Table 1 summarizes the comparative 3D tracking performance, utilizing depth and camera poses obtained from state-of-the-art 3D reconstruction models [23] for each frame. TAPIP3D consistently outperforms recent state-of-the-art methods, including SpatialTracker [28] and DELTA [44], across all three subsets in terms of AJ. We also visualize the 3D point tracking results in Figure 4; 3D point tracks from TAPIP3D are noticeably more accurate than those of the baselines. Notably, on the ADT dataset [30], characterized by particularly challenging camera motions, TAPIP3D-world significantly surpasses DELTA [28], achieving improvements of 2.5 points in AJ and 3.5 points in APD3D. Furthermore, TAPIP3D-world demonstrates a clear advantage over TAPIP3D-camera, highlighting the benefits of conducting tracking in a persistent 3D world space rather than in camera-centric coordinates.
Next, we present results for 3D point tracking on the real-world manipulation dataset DexYCB [4], as shown in Table 4. TAPIP3D notably boosts performance over the previous best method [28], improving AJ from 26.4 to 30.3. Additionally, we observe that SpaTracker [44] achieves limited performance, even falling below the depth-lifted CoTracker3 [17]. This suggests that SpaTracker’s tri-plane representation struggles to leverage the accurate sensor depth provided by DexYCB effectively.
@@
| CoTracker3 [17] + M-SaM | 19.0 | 28.1 | 75.0 | 88.0 | 88.5 | 20.0 | 30.2 | 61.4 | 80.0 | 86.2 |
| SpatialTracker [44] + M-SaM | 14.8 | 22.8 | 73.8 | 88.5 | 85.7 | 6.6 | 12.0 | 54.1 | 72.8 | 81.7 |
| DELTA [28] + M-SaM | 18.9 | 28.0 | 75.2 | 90.2 | 83.7 | 18.0 | 27.2 | 61.3 | 77.3 | 81.0 |
[…]
@@
| CoTracker3 [17] + GT | 28.4 | 35.0 | 75.0 | 88.0 | 88.5 | 27.4 | 38.0 | 61.4 | 80.0 | 86.2 |
| SpatialTracker [44] + GT | 8.2 | 13.3 | 70.6 | 85.9 | 84.9 | 7.6 | 13.6 | 56.4 | 74.7 | 83.8 |
| DELTA [28] + GT | 37.7 | 50.1 | 72.4 | 88.4 | 82.3 | 27.4 | 37.7 | 65.6 | 80.5 | 83.8 |
[…]
@@
Lastly, we evaluate our model’s 3D point tracking performance on the synthetic LSFOdyssey [37] and DynamicReplica [16] benchmarks, as shown in Table 2. Comparisons are conducted between our method and baselines under different depth sources: ground-truth depth from simulation, and estimated depth from MegaSAM [23]. We report two notable findings: 1) TAPIP3D-world outperforms all competing 3D point trackers in AJ when provided with depth estimation from either MegaSAM [23] or ground-truth depth. Moreover, utilizing ground-truth depth, our approach even outperforms the 2D tracking accuracy of CoTracker3 [17]. 2) As the depth quality improves (MegaSAM [23] GT Depth), the performance enhancement of our method in APD is substantially greater compared to DELTA [28] and SpaTracker [44], underscoring the effectiveness of our 3D point feature representation in world coordinates.

### v3

@@ 1 Introduction
Tracking points over time in video—especially through occlusions—has become a valuable tool in robotics and action recognition [cotracker, cotracker3, harley2022particle]. Fine-grained motion estimation at the particle level offers a unified framework for capturing temporal changes in object pose, part articulation, deformable structures, and even granular materials. However, most existing point trackers operate either directly in pixel space [cotracker, cotracker3, harley2022particle] or in pixel space augmented with depth information [SpatialTracker, delta]. Yet, the predominant source of apparent motion in video is often camera movement, not object motion. Since real-world dynamics unfold in 3D, tracking points directly in 3D space—rather than in the image plane—may be more natural and effective.
@@
We test TAPIP3D in established 3D point tracking benchmarks of TAPVid3D [tapvid3d], LSFOdyssey [scenetracker], Dynamic Replica [karaev2023dynamicstereo] and DexYCB [chao2021dexycb] for tracking points in 2D and 3D, using depth from ground-truth, sensors, or estimators, in camera and world 3D coordinate frames. We show TAPIP3D outperforms all previous methods in 3D point tracking metrics, with an especially large margin when accurate depth is available. While existing methods pursue 3D tracking in camera coordinates [delta, SpatialTracker], our experiments show that it is better to track in a “world” coordinate frame, which can be estimated by recent camera and depth estimation methods such as MegaSaM [megasam]. In addition to showing the impact of coordinate system choices, we ablate the contributions to performance from 3D-centric featurization and quality of depth estimation.
In summary, we present TAPIP3D for 3D point tracking from RGB-D and RGB videos that uses 3D feature clouds for video to estimate 3D point tracks, which demonstrates state-of-the-art performance when ground-truth / sensor depth is available and competitive performance with estimated depth [megasam]. It can exploit recent advances in depth and camera pose estimation [megasam, wen2025foundationstereozeroshotstereomatching] and deliver both camera and world-centric 3D tracks. To our knowledge, TAPIP3D is a pioneering 3D tracking method capable of tracking in a 3D world-centric space, with camera motion factored out.
@@ 2 Related Work
Recent advances in point tracking formulate the task as the estimation of multi-frame point trajectories, moving beyond traditional pairwise optical flow estimation. Inspired by earlier work on particle video [sand2008particle], PIPs [harley2022particle] introduces an iterative refinement approach with dense cost maps. This work along with the Tracking Any Point (TAP) benchmark [tapvid] have catalyzed a series of advancements [li2024taptr, li2024taptrv2] including the use of multi-point context [cotracker, cotracker3], enhanced initialization schemes [tapir], broader correlation contexts [bian2023context, locotrack], and post-hoc densification techniques [le2024dense].
Beyond 2D, recent works such as SpatialTracker [SpatialTracker] have extended point tracking to 3D by incorporating depth information. SpatialTracker employs a Triplane representation [triplane], where 3D featured point clouds per frame from pixel (UV) and depth (D) coordinates are projected onto onto three orthogonal planes. Triplane featurization is faster than our -NN attention, but this speedup comes at the cost of decreased performance. The works of DELTA [delta] and SceneTracker [scenetracker] also operate on a UVD coordinate system to separately compute appearance and depth correlation features, directly extending the 2D method of CoTracker [cotracker]. DELTA [delta] proposes a coarse-to-fine trajectory estimation that manages to estimate dense point tracks across the whole image plane, instead of on a set of sparse locations. In contrast to previous 3D point trackers, TAPIP3D leverages an explicit spatio-temporal 3D feature cloud video representation, instead of a 2.5D one of the image-space and depth, for feature extraction and tracking. By fully utilizing the underlying geometric structure of the scene, our method achieves better point tracking accuracy and consistency. Moreover, our model can estimate trajectories in a 3D world space, by canceling inferred camera motion from the scene image pixels during lifting, which helps performance further.
@@
3D point tracks can also be extracted from monocular videos through test-time optimization by fitting dynamic radiance fields, such as dynamic NeRFs [park2021nerfies, wang2023tracking] or Gaussians [luiten2023dynamic, wang2024shape, seidenschwarz2024dynomo, dreamscene4d] to the observed video frames supervised from reprojection RGB, depth and motion error. These approaches require per-video optimization and are thus computationally expensive. In contrast, TAPIP3D is a learning-based feed-forward approach for 3D point tracking.
@@
Scene flow estimation extends optical flow into three dimensions by considering pairs of point clouds and estimating the 3D motion between them [hadfield2011kinecting, hornacek2014sphereflow, quiroga2014dense, teed2021raft, liu2019flownet3d, wang2020flownet3d, gu2019hplflownet, niemeyer2019occupancy]. Recent works incorporate rigid motion priors, explicitly [teed2021raft] or implicitly [yang2021learning] or leveraging diffusion models [liu2024difflow3drobustuncertaintyawarescene]. Linking 3D scene flow estimates terminates at pixel occlusions by design, similarly to linking 2D flow estimates. Instead our work focuses on inferring multiframe point trajectories through occlusions.
@@
Recent methods in learning based camera motion and depth estimation, such as MoGe [moge], DUSt3R [dust3r], MonST3R [monst3r], and MegaSaM [megasam], focus on predicting dense 3D reconstructions from single images, image pairs, or videos, bringing the vision of video to 4D translation closer. MegaSaM [megasam] delivers highly accurate camera motion and depth by combining learning based initialization and updates with second order optimization, extending earlier work of DROID-SLAM [teed2021droid]. None of the methods above addresses the problem of estimating the 3D point motion, which is the focus of our work. TAPIP3D builds upon the progress of recent learning-based 3D camera motion and depth estimation methods to explore their use in 3D point tracking, by lifting videos to a 3D world space which makes the scene points easier to track than in the original 2D image plane.
@@ 3 Method
Concretely, we take as input an RGB video of frames , a sequence of depth maps , camera intrinsics (constant over time), and optionally, a sequence of camera poses. The input depth maps can originate from an off-the-shelf monocular depth estimator [moge], depth sensors, or GT depth provided by simulation environments. Camera intrinsics and extrinsics can either be estimated algorithmically by video 3D reconstruction models [megasam] or be provided by the dataset. We are given a set of different query points , , where represents the timestep at which a query is provided. Queries may also be specified in 2D, in which case we lift them to 3D using depth and camera intrinsics before sending them to the model. For each query, our model produces a 3D point trajectory , and a visibility trajectory . The architecture of TAPIP3D is illustrated in Figure 2.
@@
Previous 3D point trackers [delta, scenetracker, SpatialTracker] represent 3D point trajectories by concatenating depth values to pixel coordinates, which we refer to as space. In contrast, we use coordinates, which allows us to better leverage 3D priors, and also compensate for camera motion by transforming all frames’ points into the first camera’s coordinate system (if poses are available). Disentangling camera motion from scene motion is helpful because it makes the 3D tracks more predictable.
@@ 3.1 Preliminaries: Point tracking as iterative trajectory refinement
TAPIP3D builds on the trend of iterative architectures for point tracking [harley2022particle, tapir, cotracker, cotracker3]. We build from the most recent such tracker, CoTracker3 [cotracker3].
@@
CoTracker3 tracks 2D points across time by iteratively refining their coordinates and visibilities through a transformer. The transformer model is designed to operate on fixed-length sequences of frames only, and longer videos are processed in a sliding-window manner. Given such a sequence and a set of query points, for each query it maintains and updates per-frame (2D) locations , and confidences , and visibilities . The trajectory is initialized by copying the query coordinate to all frames (zero motion assumption); and are initialized with zeros. The model then performs iterations of refinement, producing additive updates to the coordinates and visibilities and confidences.
@@ 3.2 Tracking in a Persistent 3D Geometric Space
Our key innovation is to represent the input video as a sequence of multi-scale 2D feature maps enriched with 3D coordinates [jain2024odin, jain2025unifying]: each 2D cell holds both a -dimensional feature vector and a 3D position, where the 3D positions come from unprojected depths. This representation allows us to treat the data as a sequence of featurized point clouds (or “feature clouds”), despite their 2D grid arrangement in memory. Compared to other 3D scene representations such as tri-plane encoders [SpatialTracker] and UVD coordinates [delta], our representation encodes 3D geometry more explicitly and without distortions or approximations. We first encode each video frame with a 2D image encoder with downsampling rate , producing a sequence of feature maps at a reduced resolution , where and are the original width and height of the image. We convert these into levels of multi-scale features via 2D downsampling: for . Our downsampling uses nearest-neighbor interpolation for coordinates and average pooling for features.
@@
Our next step is to contextualize each timestep of each trajectory () with information local to it in the input data. Previous methods compute correlation features for each trajectory point using 2D neighborhoods, either on the image plane [cotracker, locotrack] or on triplanes [SpatialTracker]. Our representation enables us to use 3D neighborhoods at this step, which lets the model take advantage of the fact that the targets are in a 3D world and are often disambiguated by 3D distance. Our overall strategy is illustrated in Figure 3.
@@
These trajectory tokens are processed by a transformer utilizing proxy tokens, following CoTracker3 [cotracker3]. That is, we use virtual trajectories (that do no correspond to any query points) to help approximate spatial and temporal attention across our main tokens. The transformer outputs additive updates for positions and visibility scores:
@@
Our model is trained on the Kubric MOVi-F dataset [Greff_2022_CVPR]. We initialize the image encoder with CoTracker3’s pre-trained weights [cotracker3]. We train on 8 L40S GPUs with a batch size of 1 for 200K iterations, which takes roughly 4.2 days and consumes only 20GB VRAM with BF16 mixed precision. We train and test with a window length of . We train with videos of length 24 (making two windows per batch), and trajectories per sample. We scale the coordinates in the point maps to unit variance before processing each window to simplify learning. We optimize using AdamW [loshchilov2019decoupledweightdecayregularization] with the learning rate and weight decay both set to 5e-4. We use a learning rate schedule with 10k warmup steps followed by cosine annealing. During inference, under BF16 mixed precision, our model achieves a speed of 10 FPS and consumes around 2.6GB of VRAM when tracking 1024 query points across 32 frames on an L40S GPU.
@@ 4 Experiments
We adopt the 3D point tracking metrics from TAPVid-3D [tapvid3d]: APD3D () quantifies the average percentage of visible points whose 3D positional errors fall within depth-adaptive thresholds that scale with the ground-truth depth. Occlusion Accuracy (OA) measures the correctness of visibility predictions, while Average Jaccard (AJ) jointly evaluates position and visibility accuracy through the intersection-over-union between correctly predicted visible points and ground truth.
@@
| CoTracker3 [cotracker3] + M-SaM | 20.4 | 30.1 | 89.8 | 14.1 | 20.3 | 88.5 | 17.4 | 27.2 | 85.0 | 17.3 | 25.9 | 87.8 |
| SpatialTracker [SpatialTracker] + M-SaM | 15.9 | 23.8 | 90.1 | 7.7 | 13.5 | 85.2 | 15.3 | 25.2 | 78.1 | 13.0 | 20.8 | 84.5 |
| DELTA [delta] + M-SaM | 21.0 | 29.3 | 89.7 | 14.6 | 22.2 | 88.1 | 17.7 | 27.3 | 81.4 | 17.8 | 26.3 | 86.4 |
@@
We evaluate the performance of our model and baselines in the following synthetic and real world datasets: 1) TAPVid3D [tapvid3d] integrates videos from three distinct real-world datasets covering diverse scenarios: DriveTrack [balasingam2024drivetrack], Panoptic Studio [joo2015panoptic], and Aria Digital Twin [pan2023aria]. Together, these datasets provide a total of 4,569 evaluation videos, with video lengths ranging from 25 to 300 frames. 2) DexYCB-Pt: a dataset we introduce by leveraging ground-truth object and hand poses from manipulation scenes in DexYCB [chao2021dexycb] to generate accurate 3D point tracks on 8,000 real-world RGB-D videos. 3) LSFOdyssey [scenetracker]: a synthetic dataset that contains 90 videos with complex but realistic long-range motion using humanoids, robots, and animals. Each video is 40 frames long and has 3D point trajectories and occlusion labels, sourced from PointOdyssey [zheng2023point]. 4) DynamicReplica [karaev2023dynamicstereo]: a synthetic 3D dataset that contains 500 videos of articulated human and animal models. Each video is 300 frames long.
@@
We compare TAPIP3D against the following 2D and 3D point trackers: 1) DELTA [delta]: a very recent state-of-the-art 3D point tracker that uses depth as an additional input for computing cross-correlations, essentially extending CoTracker [cotracker] to a UVD space. 2) SpatialTracker [SpatialTracker]: a 3D point tracker that builds upon CoTracker’s architecture and uses a tri-plane representation to featurize coordinates. 3) CoTracker3 [cotracker3], which operates in 2D, but we lift its trajectories to 3D using the depth at the trajectory coordinates.
@@
In all the evaluations, our model is trained on the Kubric MOVi-F dataset [Greff_2022_CVPR]. TAPIP3D-world and TAPIP3D-camera adopt the same trained checkpoint but conduct inference in the two different 3D coordinate systems.
@@ 4.1 Point Tracking Evaluation
We first evaluate TAPIP3D against strong baseline methods [cotracker3, SpatialTracker, delta] on the challenging real-world TAPVid-3D [tapvid3d] benchmark. Table 1 summarizes the comparative 3D tracking performance, utilizing depth and camera poses obtained from state-of-the-art 3D reconstruction models [megasam] for each frame. TAPIP3D consistently outperforms recent state-of-the-art methods, including SpatialTracker [delta] and DELTA [SpatialTracker], across all three subsets in terms of AJ. We also visualize the 3D point tracking results in Figure 4; 3D point tracks from TAPIP3D are noticeably more accurate than those of the baselines. Notably, on the ADT dataset [pan2023aria], characterized by particularly challenging camera motions, TAPIP3D-world significantly surpasses DELTA [delta], achieving improvements of 2.5 points in AJ and 3.5 points in APD3D. Furthermore, TAPIP3D-world demonstrates a clear advantage over TAPIP3D-camera, highlighting the benefits of conducting tracking in a persistent 3D world space rather than in camera-centric coordinates.
Next, we present results for 3D point tracking on the real-world manipulation dataset DexYCB [chao2021dexycb], as shown in Table 4. TAPIP3D notably boosts performance over the previous best method [delta], improving AJ from 26.4 to 30.3. Additionally, we observe that SpaTracker [SpatialTracker] achieves limited performance, even falling below the depth-lifted CoTracker3 [cotracker3]. This suggests that SpaTracker’s tri-plane representation struggles to leverage the accurate sensor depth provided by DexYCB effectively.
@@
| CoTracker3 [cotracker3] + M-SaM | 19.0 | 28.1 | 75.0 | 88.0 | 88.5 | 20.0 | 30.2 | 61.4 | 80.0 | 86.2 |
| SpatialTracker [SpatialTracker] + M-SaM | 14.8 | 22.8 | 73.8 | 88.5 | 85.7 | 6.6 | 12.0 | 54.1 | 72.8 | 81.7 |
| DELTA [delta] + M-SaM | 18.9 | 28.0 | 75.2 | 90.2 | 83.7 | 18.0 | 27.2 | 61.3 | 77.3 | 81.0 |
| [2.0pt/2pt] TAPIP3D-camera + M-SaM | 20.2 | 28.6 | 70.4 | 86.2 | 83.8 | 20.6 | 30.3 | 55.3 | 77.1 | 80.2 |
@@
| CoTracker3 [cotracker3] + GT | 28.4 | 35.0 | 75.0 | 88.0 | 88.5 | 27.4 | 38.0 | 61.4 | 80.0 | 86.2 |
| SpatialTracker [SpatialTracker] + GT | 8.2 | 13.3 | 70.6 | 85.9 | 84.9 | 7.6 | 13.6 | 56.4 | 74.7 | 83.8 |
| DELTA [delta] + GT | 37.7 | 50.1 | 72.4 | 88.4 | 82.3 | 27.4 | 37.7 | 65.6 | 80.5 | 83.8 |
| [2.0pt/2pt] TAPIP3D-camera + GT | 68.3 | 83.2 | 76.0 | 91.2 | 86.2 | 53.7 | 70.8 | 64.6 | 84.7 | 84.7 |
@@
Lastly, we evaluate our model’s 3D point tracking performance on the synthetic LSFOdyssey [scenetracker] and DynamicReplica [karaev2023dynamicstereo] benchmarks, as shown in Table 2. Comparisons are conducted between our method and baselines under different depth sources: ground-truth depth from simulation, and estimated depth from MegaSaM [megasam]. We report two notable findings: 1) TAPIP3D-world outperforms all competing 3D point trackers in AJ when provided with depth estimation from either MegaSaM [megasam] or ground-truth depth. Moreover, utilizing ground-truth depth, our approach even outperforms the 2D tracking accuracy of CoTracker3 [cotracker3]. 2) As the depth quality improves (MegaSaM [megasam] GT Depth), the performance enhancement of our method in APD is substantially greater compared to DELTA [delta] and SpaTracker [SpatialTracker], underscoring the effectiveness of our 3D point feature representation in world coordinates.
| Methods | DexYCB-Pt | ||
| AJ | APD | OA | |
| CoTracker3 + SD [cotracker3] | 14.9 | 26.1 | 70.9 |
| SpatialTracker [SpatialTracker] | 5.5 | 11.4 | 66.8 |
| DELTA [delta] | 26.4 | 43.3 | 72.8 |
| TAPIP3D (Ours) | 30.3 | 52.4 | 71.3 |